The volume of a cube is increasing at a rate of 1200 cm3/min at the

moment the lengths of the sides are 20 cm. How fast are the lengths of

the sides increasing at that moment?

[8]


1
Expert's answer
2022-04-22T11:16:12-0400

"V=a^3."

"\frac{dV}{dt}=3a^2\frac{da}{dt}."

"\frac{da}{dt}=\frac{1}{3a^2}\frac{dV}{dt}=\frac{1200}{3*20^2}=1\frac{cm}{min}."